### v1

@@ I Introduction
The model is widely used for linear dimensionality reduction due to the typical situation that $r\ll p$. Suppose that we have obtained $N$ i.i.d. samples $\mathbf{y}_{i}$ from the model. The problem is about estimating the loading matrix $\bm{\Gamma}$, or equivalently, the rank $r$ symmetric matrix $\hat{\mathbf{L}}:=\bm{\Gamma}\bm{\Gamma}^{\top}\in\mathbb{R}^{p\times p}$, and the noise covariance matrix $\hat{\mathbf{S}}$. From the viewpoint of system identification, this is a static identification problem in which the output $\mathbf{y}_{i}$ depends solely on the current input $\mathbf{u}_{i}$.
@@ II Problem Formulation and Existence of a Solution
$\underset{\bm{\Sigma},\mathbf{L},\mathbf{S}}{\operatorname{min}}$ $\operatorname{tr}(\mathbf{L})+C\|\mathbf{S}\|_{0}+\mu\mathcal{D}_{\mathrm{KL}}(\mathbf{\Sigma}\|\hat{\mathbf{\Sigma}})$ (4a)
@@
In comparison with the problem formulation in [15], we do not impose the diagonal constraint on the sparse structure of $\mathbf{S}$ in (4). In this respect, our formulation is expected to be more flexible and realistic.
@@
$\operatorname{tr}(\bm{\Sigma}^{k}\check{\bm{\Sigma}}^{-1})-\log\det\bm{\Sigma}^{k}$
$=$ $\operatorname{tr}(\mathbf{X}\bm{\Lambda}^{k}\mathbf{X}^{\top})-\log\det\bm{\Sigma}^{k}$ (11a)
$=$ $\sum_{i=1}^{p}(c_{i}\lambda_{k,i}-\log\lambda_{k,i}),$ (11b)
where $\mathbf{X}=\hat{\bm{\Lambda}}^{-1/2}\hat{\mathbf{U}}^{\top}\mathbf{U}^{k}$, and $c_{i}$ is the norm squared of the $i$-th column of $\mathbf{X}$. Clearly, each $c_{i}$ is bounded away from zero because of the structure of $\mathbf{X}$. Consider the function $cx-\log x$ defined for $x>0$ with a parameter $c>0$. It is strictly convex on the positive semiaxis and has a minimum value of $1+\log c$ at $x=1/c$. Consequently, each summand in (11b) is bounded from below and $c_{1}\lambda_{k,1}-\log\lambda_{k,1}\to\infty$ because $\lambda_{k,1}\to\infty$ as argued above. Therefore, $f(\mathbf{L}^{k},\mathbf{S}^{k})$ must also diverge in this case. ∎
@@ III-C P-stationarity conditions
    A global minimizer $(\mathbf{L}^{*},\mathbf{S}^{*})$ of (5) is a P-stationary point for any $0<\gamma<{1}/{(2K)}$ if the sparse component $\mathbf{S}^{*}$ is positive definite.
@@
    Any P-stationary point is a local minimizer of (5).
@@
To show Point 1), we first fix $\mathbf{S}^{*}\succ 0$, and (20b) is just a feasibility condition. Obviously, we have
@@
where the right hand side is a smooth convex optimization problem. Hence the usual KKT conditions read as (20a), $\bm{\Lambda}^{*}\succeq 0$ in (20c), $\operatorname{tr}(\bm{\Lambda}^{*}\mathbf{L}^{*})=0$ in (20d), and (20e), where $\bm{\Lambda}^{*}$ is the Lagrange multiplier.
Then we fix $\mathbf{L}^{*}$ and derive the remaining conditions, namely $\bm{\Theta}^{*}\succeq 0$ in (20c), $\operatorname{tr}(\bm{\Theta}^{*}\mathbf{S}^{*})=0$ in (20d), and (20f). Similarly, we have
$\mathbf{S}^{*}=\underset{\mathbf{S}\succeq 0,\,\mathbf{L}^{*}+\mathbf{S}\succ 0}{\operatorname{argmin}}\ f(\mathbf{L}^{*},\mathbf{S})+C\|\mathbf{S}\|_{0}$ (22)
@@
which is to be shown next. By global optimality, we have
$f(\mathbf{L}^{*},\mathbf{S}^{*})+C\|\mathbf{S}^{*}\|_{0}\leq f(\mathbf{L}^{*},\mathbf{Z})+C\|\mathbf{Z}\|_{0}$ (24)
where $\mathbf{Z}$ is an arbitrary positive semidefinite matrix that satisfies $\mathbf{L}^{*}+\mathbf{Z}\succ 0$. Then using the mean value theorem, Cauchy–Schwarz inequality, and Assumption 2, it is not difficult to show that
@@
$=f(\mathbf{L}^{*},\mathbf{S}^{*})+\langle{\bm{\Theta}_{1}^{*}},\,{\mathbf{Z}-\mathbf{S}^{*}}\rangle$
$\quad+\langle{\nabla_{\mathbf{S}}f(\mathbf{L}^{*},\bar{\mathbf{S}})-\bm{\Theta}_{1}^{*}},\,{\mathbf{Z}-\mathbf{S}^{*}}\rangle$
$\leq f(\mathbf{L}^{*},\mathbf{S}^{*})+\langle{\bm{\Theta}_{1}^{*}},\,{\mathbf{Z}-\mathbf{S}^{*}}\rangle+K\|\mathbf{Z}-\mathbf{S}^{*}\|_{\mathrm{F}}^{2}$
where $\bar{\mathbf{S}}=t\mathbf{Z}+(1-t)\mathbf{S}^{*}$ for some $10<t<$ 1, and we have set $\bm{\Theta}_{1}^{*}:=\nabla_{\mathbf{S}}f(\mathbf{L}^{*},\mathbf{S}^{*})$ for short.
Now, let us take a particular $\mathbf{Z}=\operatorname{prox}_{\gamma C\|\cdot\|_{0}}(\mathbf{S}^{*}-\gamma\bm{\Theta}_{1}^{*})$ which is the left side of (23) with $0<\gamma<{1}/{(2K)}$. According to the definition of the proximal operator (16), we have
$C\|\mathbf{Z}\|_{0}+\frac{1}{2\gamma}\|\mathbf{Z}-(\mathbf{S}^{*}-\gamma\bm{\Theta}_{1}^{*})\|_{\mathrm{F}}^{2}\leq C\|\mathbf{S}^{*}\|_{0}+\frac{\gamma}{2}\|\bm{\Theta}_{1}^{*}\|_{\mathrm{F}}^{2}.$ (26)
@@
$\leq C\|\mathbf{Z}\|_{0}-C\|\mathbf{S}^{*}\|_{0}+\langle{\bm{\Theta}_{1}^{*}},\,{\mathbf{Z}-\mathbf{S}^{*}}\rangle+K\|\mathbf{Z}-\mathbf{S}^{*}\|_{\mathrm{F}}^{2}$ (27b)
$=C\|\mathbf{Z}\|_{0}-C\|\mathbf{S}^{*}\|_{0}+\frac{1}{2\gamma}\|\mathbf{Z}-(\mathbf{S}^{*}-\gamma\bm{\Theta}_{1}^{*})\|_{\mathrm{F}}^{2}$
$\quad-\frac{\gamma}{2}\|\bm{\Theta}_{1}^{*}\|_{\mathrm{F}}^{2}+(K-\frac{1}{2\gamma})\|\mathbf{Z}-\mathbf{S}^{*}\|_{\mathrm{F}}^{2}$
$\leq(K-\frac{1}{2\gamma})\|\mathbf{Z}-\mathbf{S}^{*}\|_{\mathrm{F}}^{2}\leq 0$ (27c)
@@
Now let us treat Point 2). Suppose that we are given a P-stationary point in the sense of Definition 1. For simplicity, we write $\bm{\Phi}^{*}=(\mathbf{L}^{*},\mathbf{S}^{*})$ and still $\bm{\Theta}_{1}^{*}=\nabla_{\mathbf{S}}f(\mathbf{L}^{*},\mathbf{S}^{*})$. The aim is to show that there exists a sufficiently small neighborhood
@@
with $\delta>0$ such that $\bm{\Phi}^{*}$ is locally optimal in $U(\bm{\Phi}^{*},\delta)\cap\mathscr{D}$ where $\mathscr{D}$ is the feasible set given in (6). That is to say, for all $(\mathbf{L},\mathbf{S})\in U(\bm{\Phi}^{*},\delta)\cap\mathscr{D}$, it holds that
@@
We take $\delta=\min\{\delta_{1},\delta_{2}\}$ in (28). In order to show the inequality (29), we need to introduce some index sets. Let $\mathbb{N}_{p}:=\{1,2,\dots,p\}$ and $\mathbb{N}_{p}^{2}=\mathbb{N}_{p}\times\mathbb{N}_{p}$ be the Cartesian product providing indices for the entries of $p\times p$ matrices. Define $\Gamma_{*}:=\{(i,j)\in\mathbb{N}_{p}^{2}:s^{*}_{ij}=0\}$ and $\overline{\Gamma}_{*}:=\mathbb{N}_{p}^{2}\backslash\Gamma_{*}$ the complement. According to the evaluation formula (15) for the proximal operator, we have the following observations.
  • For $(i,j)\in\Gamma_{*}$, the elementwise version of (20f) reads as $\operatorname{prox}_{\gamma C|\cdot|_{0}}[s^{*}_{ij}-\gamma(\theta^{*}_{1,ij}-\theta^{*}_{ij})]=s^{*}_{ij}=0$. This implies that
    $|\gamma(\theta^{*}_{1,ij}-\theta^{*}_{ij})|\leq\sqrt{2\gamma C}\iff|\theta_{1,ij}^{*}-\theta^{*}_{ij}|\leq\sqrt{2C/\gamma}.$ (32)
  • For $(i,j)\in\overline{\Gamma}_{*}$ on the contrary, we have $s_{ij}^{*}\neq 0$ and $|s^{*}_{ij}-\gamma(\theta^{*}_{1,ij}-\theta^{*}_{ij})|>\sqrt{2\gamma C}$, which in view of (15), leads to
    $s^{*}_{ij}-\gamma(\theta^{*}_{1,ij}-\theta^{*}_{ij})=s_{ij}^{*}\iff\theta_{1,ij}^{*}=\theta^{*}_{ij}.$ (33)
@@
=\,= $\langle{\bm{\Lambda}^{*}},\,{\mathbf{L}-\mathbf{L}^{*}}\rangle+\langle{\bm{\Theta}_{1}^{*}},\,{\mathbf{S}-\mathbf{S}^{*}}\rangle$ (35b)
=\,= $\langle{\bm{\Lambda}^{*}},\,{\mathbf{L}}\rangle+\langle{\bm{\Theta}_{1}^{*}},\,{\mathbf{S}-\mathbf{S}^{*}}\rangle$ (35c)
@@
=\,= $\sum_{(i,j)\in\Gamma_{*}}(\theta_{1,ij}^{*}-\theta_{ij}^{*})(s_{ij}-s^{*}_{ij})$ (35e)
=\,= $\sum_{(i,j)\in\Gamma_{*}}(\theta_{1,ij}^{*}-\theta_{ij}^{*})s_{ij}=0,$ (35f)
@@
  • (35b) results from substitutions using (20e) and the definition of $\bm{\Theta}_{1}$,

### v2

@@ I Introduction
The model is widely used for linear dimensionality reduction due to the typical situation that $r\ll p$. Suppose that we have obtained $N$ i.i.d. samples $\mathbf{y}_{i}$ from the model. The problem is about estimating the symmetric matrix $\hat{\mathbf{L}}:=\bm{\Gamma}\bm{\Gamma}^{\top}\in\mathbb{R}^{p\times p}$ of rank $r$ and the noise covariance matrix $\hat{\mathbf{S}}$. From the viewpoint of system identification, this is a static identification problem in which the output $\mathbf{y}_{i}$ depends solely on the current input $\mathbf{u}_{i}$.
@@ II Problem Formulation and Existence of a Solution
$\underset{\bm{\Sigma},\mathbf{L},\mathbf{S}}{\operatorname{min}}$ $\operatorname{tr}(\mathbf{L})+C\|\mathbf{S}\|_{0}+\mu\mathcal{D}_{\mathrm{KL}}(\mathbf{\Sigma}\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\check{\mathbf{\Sigma}}})$ (4a)
@@
The problem formulation in [15] imposes the diagonal constraint on the sparse structure of $\mathbf{S}$ which can be understood as prior information about the noise. In the absence of such prior information, our formulation (4), which only requires $\mathbf{S}$ to be sparse, is expected to be more flexible and realistic.
@@
$\operatorname{tr}(\bm{\Sigma}^{k}\check{\bm{\Sigma}}^{-1})-\log\det\bm{\Sigma}^{k}$ $=\operatorname{tr}(\mathbf{X}\bm{\Lambda}^{k}\mathbf{X}^{\top})-\log\det\bm{\Sigma}^{k}$
$=\sum_{i=1}^{p}(c_{i}\lambda_{k,i}-\log\lambda_{k,i}),$ (11a)
where $\mathbf{X}=\hat{\bm{\Lambda}}^{-1/2}\hat{\mathbf{U}}^{\top}\mathbf{U}^{k}$, and $c_{i}$ is the norm squared of the $i$-th column of $\mathbf{X}$. Clearly, each $c_{i}$ is bounded away from zero because of the structure of $\mathbf{X}$. Consider the function $cx-\log x$ defined for $x>0$ with a parameter $c>0$. It is strictly convex on the positive semiaxis and has a minimum value of $1+\log c$ at $x=1/c$. Consequently, each summand in (11a) is bounded from below and $c_{1}\lambda_{k,1}-\log\lambda_{k,1}\to\infty$ because $\lambda_{k,1}\to\infty$ as argued above. Therefore, $f(\mathbf{L}^{k},\mathbf{S}^{k})$ must also diverge in this case. ∎
@@ III-C P-stationarity conditions
    If $(\mathbf{L}^{*},\mathbf{S}^{*})$ is a global minimizer of (5) and the sparse component $\mathbf{S}^{*}$ is positive definite, then $(\mathbf{L}^{*},\mathbf{S}^{*})$ is a P-stationary point where the parameter $\gamma$ can take an arbitrary value in the interval $(0,{1}/{(2K)})$.
@@
    If $(\mathbf{L}^{*},\mathbf{S}^{*})\in\mathscr{D}$ is a P-stationary point, then $(\mathbf{L}^{*},\mathbf{S}^{*})$ is a local minimizer of (5).
@@
To show Point 1), let us suppose that $(\mathbf{L}^{*},\mathbf{S}^{*})$ is a global minimizer of (5), and then we need to show that the P-stationary point conditions (20) hold for $(\mathbf{L}^{*},\mathbf{S}^{*})$.
The conditions (20a) and (20b) are just feasibility. Then, we first fix $\mathbf{S}^{*}\succ 0$, and consider the optimality with respect to $\mathbf{L}$. Indeed, the term $C\|\mathbf{S}\|_{0}$ in (5) now becomes a constant $C\|\mathbf{S}^{*}\|_{0}$, which implies that
@@
where the right hand side is a smooth convex optimization problem. Hence the usual KKT conditions read as $\bm{\Lambda}^{*}\succeq 0$ in (20c), $\operatorname{tr}(\bm{\Lambda}^{*}\mathbf{L}^{*})=0$ in (20d), and (20e), where $\bm{\Lambda}^{*}$ is the Lagrange multiplier associated to the constraint $\mathbf{L}\succeq 0$.
Next, we fix $\mathbf{L}^{*}$ and derive the remaining conditions, namely $\bm{\Theta}^{*}\succeq 0$ in (20c), $\operatorname{tr}(\bm{\Theta}^{*}\mathbf{S}^{*})=0$ in (20d), and (20f) by considering the optimality with respect to $\mathbf{S}$. Similar to the characterization (21), we have
$\mathbf{S}^{*}=\underset{\begin{subarray}{c}\mathbf{S}\succeq 0,\,\mathbf{L}^{*}+\mathbf{S}\succ 0\end{subarray}}{\operatorname{argmin}}\ f(\mathbf{L}^{*},\mathbf{S})+C\|\mathbf{S}\|_{0}.$ (22)
@@
which remains to be established in the following.
By global optimality of $(\mathbf{L}^{*},\mathbf{S}^{*})$, we have
$f(\mathbf{L}^{*},\mathbf{S}^{*})+C\|\mathbf{S}^{*}\|_{0}\leq f(\mathbf{L}^{*},\mathbf{Z})+C\|\mathbf{Z}\|_{0},$ (24)
where $\mathbf{Z}$ is an arbitrary positive semidefinite matrix that satisfies $\mathbf{L}^{*}+\mathbf{Z}\succ 0$. Let $\mathbf{\Xi}^{*}:=\nabla_{\mathbf{S}}f(\mathbf{L}^{*},\mathbf{S}^{*})$. Then, we obtain the inequalities
@@
$=f(\mathbf{L}^{*},\mathbf{S}^{*})+\langle{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{\Xi}^{*}}},\,{\mathbf{Z}-\mathbf{S}^{*}}\rangle$
$\quad+\langle{\nabla_{\mathbf{S}}f(\mathbf{L}^{*},\bar{\mathbf{S}})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{\Xi}^{*}}},\,{\mathbf{Z}-\mathbf{S}^{*}}\rangle$
$\leq f(\mathbf{L}^{*},\mathbf{S}^{*})+\langle{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{\Xi}^{*}}},\,{\mathbf{Z}-\mathbf{S}^{*}}\rangle+K\|\mathbf{Z}-\mathbf{S}^{*}\|_{\mathrm{F}}^{2},$
where the first equality is due to the mean value theorem with $\bar{\mathbf{S}}=t\mathbf{Z}+(1-t)\mathbf{S}^{*}$ for some $t\in(0,1)$, and the last inequality follows from Cauchy–Schwarz and Assumption 2. Now, let us take a particular $\mathbf{Z}=\operatorname{prox}_{\gamma C\|\cdot\|_{0}}(\mathbf{S}^{*}-\gamma{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{\Xi}^{*}})$ which is the left side of (23) with $0<\gamma<{1}/{(2K)}$. According to the definition of the proximal operator (16), we have
$C\|\mathbf{Z}\|_{0}+\frac{1}{2\gamma}\|\mathbf{Z}-(\mathbf{S}^{*}-\gamma{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{\Xi}^{*}})\|_{\mathrm{F}}^{2}\leq C\|\mathbf{S}^{*}\|_{0}+\frac{\gamma}{2}\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{\Xi}^{*}}\|_{\mathrm{F}}^{2}.$ (26)
@@
$\leq C\|\mathbf{Z}\|_{0}-C\|\mathbf{S}^{*}\|_{0}+\langle{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{\Xi}^{*}}},\,{\mathbf{Z}-\mathbf{S}^{*}}\rangle+K\|\mathbf{Z}-\mathbf{S}^{*}\|_{\mathrm{F}}^{2}$ (27b)
$=C\|\mathbf{Z}\|_{0}-C\|\mathbf{S}^{*}\|_{0}+\frac{1}{2\gamma}\|\mathbf{Z}-(\mathbf{S}^{*}-\gamma{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{\Xi}^{*}})\|_{\mathrm{F}}^{2}$
$\quad-\frac{\gamma}{2}\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{\Xi}^{*}}\|_{\mathrm{F}}^{2}+(K-\frac{1}{2\gamma})\|\mathbf{Z}-\mathbf{S}^{*}\|_{\mathrm{F}}^{2}$
$\leq(K-\frac{1}{2\gamma})\|\mathbf{Z}-\mathbf{S}^{*}\|_{\mathrm{F}}^{2}\leq 0,$ (27c)
@@
Now let us treat Point 2). Suppose that $(\mathbf{L}^{*},\mathbf{S}^{*})\in\mathscr{D}$ is a P-stationary point in the sense of Definition 1, where $\mathscr{D}$ is the feasible set (6). For simplicity, we write $\bm{\Phi}^{*}=(\mathbf{L}^{*},\mathbf{S}^{*})$ and still ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{\Xi}^{*}}=\nabla_{\mathbf{S}}f(\mathbf{L}^{*},\mathbf{S}^{*})$. The aim is to show that there exists a sufficiently small neighborhood
@@
with $\delta>0$ such that $\bm{\Phi}^{*}$ is locally optimal in $U(\bm{\Phi}^{*},\delta)\cap\mathscr{D}$. That is to say, for all $(\mathbf{L},\mathbf{S})\in U(\bm{\Phi}^{*},\delta)\cap\mathscr{D}$, it holds that
@@
We take $\delta=\min\{\delta_{1},\delta_{2}\}$ in (28), and proceed to prove the local optimality of $\bm{\Phi}^{*}$, i.e., to show the inequality (29).
Let us introduce some index sets. Let $\mathbb{N}_{p}:=\{1,2,\dots,p\}$ and $\mathbb{N}_{p}^{2}=\mathbb{N}_{p}\times\mathbb{N}_{p}$ be the Cartesian product providing indices for the entries of $p\times p$ matrices. Define $\Gamma_{*}:=\{(i,j)\in\mathbb{N}_{p}^{2}:s^{*}_{ij}=0\}$ and $\overline{\Gamma}_{*}:=\mathbb{N}_{p}^{2}\backslash\Gamma_{*}$ the complement. According to the evaluation formula (15) for the proximal operator, we have the following observations.
  • For $(i,j)\in\Gamma_{*}$, the elementwise version of (20f) reads as $\operatorname{prox}_{\gamma C|\cdot|_{0}}[s^{*}_{ij}-\gamma({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi_{ij}^{*}}-\theta^{*}_{ij})]=s^{*}_{ij}=0$. This implies that
    $|\gamma({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi_{ij}^{*}}-\theta^{*}_{ij})|\leq\sqrt{2\gamma C}\iff|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi_{ij}^{*}}-\theta^{*}_{ij}|\leq\sqrt{2C/\gamma}.$ (32)
  • For $(i,j)\in\overline{\Gamma}_{*}$ on the contrary, we have $s_{ij}^{*}\neq 0$ and $|s^{*}_{ij}-\gamma({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi_{ij}^{*}}-\theta^{*}_{ij})|>\sqrt{2\gamma C}$, which in view of (15), leads to
    $s^{*}_{ij}-\gamma({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi_{ij}^{*}}-\theta^{*}_{ij})=s_{ij}^{*}\iff{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi_{ij}^{*}}=\theta^{*}_{ij}.$ (33)
@@
=\,= $\langle{\bm{\Lambda}^{*}},\,{\mathbf{L}-\mathbf{L}^{*}}\rangle+\langle{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{\Xi}^{*}}},\,{\mathbf{S}-\mathbf{S}^{*}}\rangle$ (35b)
=\,= $\langle{\bm{\Lambda}^{*}},\,{\mathbf{L}}\rangle+\langle{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{\Xi}^{*}}},\,{\mathbf{S}-\mathbf{S}^{*}}\rangle$ (35c)
@@
=\,= $\sum_{(i,j)\in\Gamma_{*}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi_{ij}^{*}}-\theta_{ij}^{*})(s_{ij}-s^{*}_{ij})$ (35e)
=\,= $\sum_{(i,j)\in\Gamma_{*}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi_{ij}^{*}}-\theta_{ij}^{*})s_{ij}=0,$ (35f)
@@
  • (35b) results from substitutions using (20e) and the definition of $\mathbf{\Xi}^{*}$,